### v1

# PillarTrack: Redesigning Pillar-based Transformer Network for Single Object Tracking on Point Clouds

Abstract
LiDAR-based 3D single object tracking (3D SOT) is a critical issue in robotics and autonomous driving. It aims to obtain accurate 3D BBox from the search area based on similarity or motion. However, existing 3D SOT methods usually follow the point-based pipeline, where the sampling operation inevitably leads to redundant or lost information, resulting in unexpected performance. To address these issues, we propose PillarTrack, a pillar-based 3D single object tracking framework. Firstly, we transform sparse point clouds into dense pillars to preserve the local and global geometrics. Secondly, we introduce a Pyramid-type Encoding Pillar Feature Encoder (PE-PFE) design to help the feature representation of each pillar. Thirdly, we present an efficient Transformer-based backbone from the perspective of modality differences. Finally, we construct our PillarTrack tracker based above designs. Extensive experiments on the KITTI and nuScenes dataset demonstrate the superiority of our proposed method. Notably, our method achieves state-of-the-art performance on the KITTI and nuScenes dataset and enables real-time tracking speed. We hope our work could encourage the community to rethink existing 3D SOT tracker designs.
I INTRODUCTION
3D Single object tracking (3D SOT) has a diverse applications in autonomous driving and robotics. Given the initial state (appearance and location) of specific target in the first frame, 3D single object tracking aims to estimate its 3D state across subsequent frames. Existing LiDAR-based 3D single object tracking methods [1, 2, 3, 4, 5, 6, 7] generally follow the Siamese paradigm from 2D visual object tracking, aiming to achieve a trade-off between runtime and accuracy. Given the sparse and irregular input point cloud, these methods initially leverage PointNet families [8, 9] to learn point-wise discriminative representation, and then obtain point-wise similarity with a feature aggregation module. Finally, they estimate the state of specific targets based on these similarity features. SC3D [10] stands as the pioneering LiDAR-based 3D Siamese tracker, leveraging an efficient auto-encoder based on PointNet[11] for point-wise feature encoding. Subsequently, approaches such as P2B [2], 3D-SiamRPN[1], BAT [4] and PTT [3, 12] employ PointNet++ [9] to extract more effective point-wise representation and achieve superior tracking performance. However, one common issue with above point-based 3D SOT methods is the need for re-sampling the input points to a fixed number. For instance. P2B [2] requires the input points to be re-sample to 1024 for the search area and 512 for the template point cloud in order to align the network input. This re-sample operation inevitably introduces the potential for redundant or lost information, which can adversely affect performance. Additionally, the reliance on point query/retrieval in 3D space (e.g., PointNet++ [9]) for the re-sampling process may not be friendly to efficient hardware implementation.
Figure 1: Comparison with other 3D SOT methods on KITTI dataset. We categorize methods based on the Backbone architecture and report their performance on Success and Precision.
To this end, inspired by the recent advancements in pillar-based 3D object detectors [13, 14, 15] due to their real-time speed and high performance. Therefore, we convert the sparse and irregular point clouds into dense and regular pillar representations. It is noteworthy that the 3D SOT task can be regarded as a specific 3D object detection task within a local search area, with the additional template point prior information. Therefore,it is natural to leverage advanced principles from 3D point cloud detection to improve point cloud tracking tasks. Specifically, pillar-based point cloud representation possesses following advantages:(1) Pillar representation is dense and ordered, facilitating seamless integration with advanced 2D image-based techniques without much modification. (2) The compact nature of the pillar representation reduces computational overhead while maintaining a desirable trade-off between performance and speed. (3) Pillar representation is deployment-friendly, making it highly suitable for resource-limited devices like mobile robots or drones. Particularly,pillar representation naturally aligns with the high demanding real-time requirements of 3D trackers, making it well-suited for tracking tasks.
Currently, few works [16, 17] are dedicated to pillar-based 3D SOT with point clouds. SMAT [16] addresses this gap by transforming sparse 3D point clouds into dense pillars and then utilizing a transformer-based encoder for multi-scale feature extraction and interaction, yielding promising results. However, SMAT directly adopts a simplistic pillar encoding module from PointPillars [18] limits the quality of the pillar representations. Besides, their backbone directly follows the design principle of vision transformers tailored for RGB image, which may be sub-optimal for point cloud modalities. PTTR++ [17] introduces a fusion of point-wise and pillar-wise views to further enhance the performance of PTTR [6]. Nevertheless, PTTR++ still involves a re-sampling operation within its pipeline, which presents paratical deployment challenges.
In this paper, we propose PillarTrack, a pillar-based 3D single-object tracking framework, aiming to improve tracking performance while enhancing the inference speed. Firstly, we transform sparse and unordered point clouds into dense and regular pillar representations to reduce information loss caused by re-sampling operation and propose a Pyramid-type Encoding Pillar Feature Encoder (PE-PFE) design to help the feature learning of each pillar. Secondly, we argue the design of existing Transformer-based backbone from the perspective of modality differences and propose an modality-aware Transformer-based backbone specifically tailored for point cloud modalities, achieving higher performance with fewer GFLOPs. Finally, we construct our PillarTrack network. Extensive experiments on the KITTI and nuScenes dataset demonstrate the superiority of our proposed method. As shown in Figure  1, our approach achieves state-of-the-art performance on the KITTI and nuScenes dataset, striking a better balance between speed and accuracy, and catering to diverse practical needs. In conclusion, the main contributions of this paper are as follows:
$\bullet$ PE-PFE: A Pyramid-type Encoding Pillar Feature Encoder (PE-PFE) design to encode the point coordinate of each pillar with pyramid-type representation, and bring the performance gain without extra computational overhead.
$\bullet$ Modality-aware Transformer-based Backbone: A more suitable for point cloud modalities, aiming to enhance feature representation. This design involves a simple adjustment of computational resources allocated to the front-end of the backbone, allowing for capturing more semantic details of the input point cloud.
$\bullet$ SOTA and Open-source: Experiments on KITTI and nuScenes dataset show our method achieve the state-of-the-art performance. Besides, we will open source our code to the research community in https://github.com/StiphyJay/PillarTrack.
II Related Work
II-A Single Object Tracking on Point Clouds
Benefiting from the nature of LiDAR that are less sensitive to illumination changes and capture accurate distance information, numerous works[10, 19, 20, 1, 2, 4, 21, 22] in LiDAR-based 3D Single Object Tracking (3D SOT) have merged. SC3D[10] leverages the Kalman filter for proposals generation and select the most suitable candidate based on Siamese network[23]. 3D-SiamRPN [1] combines 3D Siamese network with 3D RPN network for object tracking. P2B[2] introduces target-specific features and utilizes VoteNet[24] to estimate target center. Based on the P2B pipeline, BAT[4] incorporated Box-aware information to enhance similarity feature. Besides, V2B[22] introduces Voxel-to-BEV network for pixel-wise regression of 3D targets, M2-Track [21, 25] adopts a motion-centric paradigm to address the issue of 3D SOT and achieves remarkable performance. Inspired by the success of transformers[26], several methods [3, 12, 5, 6, 27, 7, 28, 29] have merged. PTT [3, 12] proposed Point-Track-Transformer module to assign weights to crucial point cloud features. LTTR[5] proposes a Transformer-based feature fusion module to learn semantic context among different region. PTTR[6] employs an attention-based feature matching mechanism to integrate target clues in similarity. STNet[7] develops an iterative coarse-to-fine correlation network for robust correlation learning. CXTrack[29] adopts transformer to extract richer contextual information from consecutive frames, enhancing localization accuracy. Recently, SMAT[16] encodes global similarity based on attention in the BEV to alleviate the impact of point sparsity. However, their backbone still partially inherits the design principles from ResNet[30] tailored for the image modality, which may be sub-optimal for point clouds modality.
Figure 2: The architecture of our PillarTrack network.Given the template and search area, we first use PE-Pillar to extract multi-scale features respectively. Then, MAE computes similarity by attention at each feature scale and then fuses the multi-scale similarity feature.Finally, we apply the detection head on the feature fusion map to localize the target.
II-B Pillar-based Object Detection on Point Clouds
Grid-based point cloud 3D detectors[18, 31, 13, 14, 15] generally transform unordered points into regular pillar/voxel and then utilize 2D/3D CNN to learn representation. VoxelNet [32] stands as a pioneering work that leverages Voxel Feature Extractor (VFE) and 3D CNN for 3D object detection. Subsequently, SECOND [33] employs 3D sparse convolution on non-empty voxels to accelerate training and inference. Different from above approaches, PointPillars [18] innovatively converts point clouds into pillars and leverages 2D CNN to enhance detection speed. After that, CenterPoint[31] proposes a center-based head design, achieving milestone performance. Furthermore, PillarNet[13] improves the accuracy of pillar-based 3D detectors through an ”encoder-neck-head” architecture, achieving superior performance while maintaining real-time speed. Based on PillarNet[13], FastPillars[14] proposes a backbone design suitable for point cloud modality and considers deployment-friendly ability, resulting in new state-of-the-art. The success of these pillar-based 3D detectors greatly motivates us to explore the potential of simple pillar-based representation in 3D SOT tasks.
II-C Vision Transformer
Transformer[26] is first proposed in natural language processing (NLP). Subsequently, owing to its remarkable ability for global modeling, the Transformer has shown significant success in image recognition tasks. ViT[34] leverage the Transformer structure to classify images by splitting them into patches. SwinTransformer[35] proposes a shifted windows attention mechanism, achieving excellent performance across various image tasks. PVT[36] and PVTv2[37] utilize a progressive shrinking pyramid to construct a visual Transformer backbone. In the 3D Vison, Point Transformer[38] introduces a specialized Point Transformer layer tailored for processing unordered point clouds. PCT[39] utilizes 3D coordinates of points to generate position embedding and employs an attention module to extract point features from local neighborhoods. These efforts showcase outstanding performance in point cloud shape classification or part segmentation. For LiDAR-based 3D detection tasks, recent approaches[40, 41] first transform point clouds into regular pillars/voxels and then utilize dynamic sparse windows on BEV feature for feature extraction, yielding remarkable performance. In contrast, we employ the Transformer for global modeling on the pillarized feature, aiming to explore the potential of the Transformer in the object tracking task from a BEV perspective.
III Preliminary
III-A Problem Statement
In the task of the 3D Single Object Tracking (3D SOT), we initially define target point cloud (i.e., template) $\mathbf{P_{t}}=\{\mathbf{p}_{ti}=[x_{i},y_{i},z_{i},r_{i}]^{T}\in\mathbb{R}^{N\times 4}\}$, where $x_{i},y_{i},z_{i}$ denote the coordinate values of each point along the axes X, Y, Z, respectively, and $r_{i}$ represents the laser reflection intensity. Our goal is to locate the the 3D bounding box (3D BBox) of this target in the search area $\mathbf{P_{t}}=\{\mathbf{p}_{si}=[x_{j},y_{j},z_{j},r_{j}]^{T}\in\mathbb{R}^{M\times 4}\}$ frame by frame. Here, $N$ and $M$ is the number of points in target and search area. The 3D BBox is represented as $\mathbf{B}=\{\mathbf{b}=[x,y,z,h,w,l,{\theta}]^{T}\in\mathbb{R}^{1\times 7}\}$, The coordinates $x,y,z$ indicate the object’s center, while $h,w,l$ denote its size, ${\theta}_{j}$ is the object’s heading angle. Since the target’s 3D BBox is known in the first frame, we need to regress the target center and heading angle in subsequent frames. By applying the displacement and heading angle from the previous frame’s 3D BBox, we can locate the target in the current frame.
III-B Baseline Architecture Overview
As SMAT represents the first attempt at a pillar-based 3D Siamese tracker and has shown promising performance, we choose SMAT as our baseline and provide a brief overview of its structure. The core components of SMAT include: 1) a naive pillar feature encoding module derived from Pointpillars, which transforms raw point cloud data into a pseudo-image; 2) a Transformer-based backbone network that computes similarity through a multi-head attention mechanism for the multi-scale features from both the template and search branches, capturing local and global information within the point cloud; 3) a neck module that integrates features from different scales and fuses multi-scale similarity features, enhancing the model’s perception of multi-scale information; 4) a decoder that predicts the position and orientation of the target object. These carefully designed features enable SMAT to effectively handle sparse point cloud data, offering an efficient solution for 3D Single Object Tracking.
III-C Motivation
Our research endeavors to investigate a real-time, high-performance, pillar-based approach for 3D single-object tracking. To achieve this, we leverage the compact pillar representation and design a strong Transformer-based backbone that enhances pillar feature representation with reduced computational resources from the perspective of modality differences. Regarding the feature integration and similarity computation module, we follow the original design of baseline method[16], which has shown remarkable efficacy. We hope that our design principles offer a novel perspective in 3D single object tracking, encouraging the community to revisit existing methodologies.
IV Method
IV-A Pyramid-Encoded Pillar Feature Encoding Design
Figure 3: The illustration of PE-PillarVFE design. We encode the channel information of the input point cloud in a pyramid-type fashion. This Pyramid-type encoding design allows the network to optimize more effectively without the input information loss.
In pillar-based 3D detection methods[18, 13, 14], Pillar Feature Encoding (PFE) plays a crucial role, directly impacting the representation learning of BEV feature and final detection accuracy. Its significance extends to 3D single object tracking tasks as well. In SMAT, they utilize the PFE module from PointPillars [18] to extract pillar-wise feature. Here, taking PFE module as an example, we illustrate the details of their input encoding. The points in each pillar are initially concatenated along the channel dimension to form $\hat{p_{i}}=\{[x_{i},y_{i},z_{i},r_{i},{x}_{i}^{m},{y}_{i}^{m},{z}_{i}^{m},{x}_{i}^{c},{y}_{i}^{c},{z}_{i}^{c}]\in{\mathbb{R}^{N_{v}\times 10}}\}$, where $[x_{i},y_{i},z_{i}]$ is the original point coordinates in the ego frame, $[{x}_{i}^{m},{y}_{i}^{m},{z}_{i}^{m}]$ denotes the offset of $p_{i}$ from the mean of the xyz coordinates of all points within the current pillar, and $[{x}_{i}^{c},{y}_{i}^{c},{z}_{i}^{c}]$ is the offset of $p_{i}$ from the current pillar center. However, there exist significant numerical differences in the point-wise feature. For example, the numerical range of the $[x_{i},y_{i},z_{i}]$ is $[-3.2,3.2]$ (i.e., the range of search area). However, the numerical range of $[{x}_{i}^{m},{y}_{i}^{m},{z}_{i}^{m}]$ and $[{x}_{i}^{c},{y}_{i}^{c},{z}_{i}^{c}]$ is $[-0.1,0.1]$, since they are relative offset within a pillar. It can be observed that there exist significant numerical differences (almost 30 $\times$) between the point-wise feature channels in PFE. Such huge disparities in input values pose significant challenges for network optimization  [42, 43]. Therefore, we propose a Pyramid-type Encoding Pillar Feature Encoder (PE-PFE) design aimed at reducing the numerical differences between the input channels, enabling the network towards better optimization on numerically stable inputs.
TABLE I: Tracking performance comparison of different PFE designs in the Car category in KITTI.
PFE Method Success Precision
PFE (SMAT [16] [^4] The model was trained on a setup comprising four NVIDIA GeForce RTX 3090 GPUs) 68.9 79.6
PE-PFE(Ours) 70.8 (+1.9) 80.3 (+0.7)
Figure 4: The loss comparison of different PFE designs.
As shown in Figure 3, we encode the channel values of the input point cloud in a pyramid-type manner, and finally retain a residual value that cannot be evenly divided. The encoding values from each layer and the residual value are then cascaded together as point-wise features, which are then processed by the PFE module. It is important to note that this pyramid-type encoding design significantly reduces the numerical differences between the input data channels without sacrificing the information contained in the point cloud. This design helps the network achieve better optimization results. Furthermore, to demonstrate the effectiveness of our PE-PFE design compared to the naive PFE from PointPillars [18], we present the performance comparison in the Car category. As shown in Table I, our simple yet effective PE-PFE design yields 1.9 point gains in Success and 0.7 point improvement in Precision. As shown in Figure 4, the design of PE-PFE leads to lower loss values, enabling the network to find better parameters and achieve superior performance.
IV-B Modality-aware Transformer-based Backbone
Recent studies [44, 14, 45] have pointed out that the backbone designs commonly employed in image domain [46, 30, 35, 47] are sub-optimal for point cloud or depth images due to significant modality differences between RGB images and LiDAR point clouds. Particularly, backbone design in image domain tends to utilize more block layers in later stages to extract high-level semantic features for perception. For instance, ResNet-50[30] takes 6 residual blocks in the 3rd stage while the 1st stage uses 3 blocks for feature extraction. Similarly, Swin-Transformer [35] adopts 9 blocks in the 3rd stage while only 1 block is used in the 1st stage. However, we think that these design choices for distributing computations across stages were largely empirical and may not be well-suited for LiDAR point clouds. Unlike RGB images, LiDAR can capture spatial distances and shapes of objects. This means that rich and accurate geometric information of objects is already explicitly encoded in point cloud. Therefore, we assume that instead of allocating excessive computational resources to model object geometries in later stages like ResNet/Swin-Transformer in RGB images, we propose reallocating more computational resources to the early stages to effectively incorporate the geometric information carried by raw point clouds.
TABLE II: Impact of varying block quantities in the four stages of the backbone on the final results. Bold denotes the best performance.
Blocks Number GFLOPs Success Precision
$\left[3,4,6,3\right]$ (SMAT[16]44footnotemark: 4) 0.37 68.9 79.6
$\left[2,2,6,2\right]$ 0.29 25.7 34.9
$\left[1,1,9,1\right]$ 0.29 53.1 56.2
$\left[1,1,1,1\right]$ 0.14 69.5 80.1
$\left[2,1,1,1\right]$ 0.16 70.3 80.6
$\left[3,1,1,1\right]$ 0.18(-0.19) 72.1(+3.2) 82.1 (+2.5)
$\left[4,1,1,1\right]$ 0.21 68.7 78.9
$\left[1,2,1,1\right]$ 0.16 70.1 80.7
$\left[1,3,1,1\right]$ 0.18 69.7 80.6
$\left[1,4,1,1\right]$ 0.20 69.4 79.4
$\left[1,1,2,1\right]$ 0.16 68.4 78.8
$\left[1,1,3,1\right]$ 0.18 70.0 80.8
$\left[1,1,4,1\right]$ 0.20 71.0 81.8
$\left[1,1,1,2\right]$ 0.15 68.0 78.3
$\left[1,1,1,3\right]$ 0.17 68.1 78.4
$\left[1,1,1,4\right]$ 0.18 69.3 80.0
To validate the aforementioned hypothesis for transformer-based backbone, we systematically study computation allocation by adjusting the stage compute ratio in the transformer-based backbone of SMAT [16]. It is worth noting that the setting of SMAT has the same depth number as ResNet-50, with a stage compute ratio of $[3,4,6,3]$. As shown in Table II, we used the SMAT backbone as the baseline and varied the number of depth blocks from 0 to 4 with a stride of 1 in each stage. From the results, we can find that the performance is quite sensitive to the capacity of early stage 1, while being less sensitive to the later stages (2, 3, 4). This confirmed that allocating more computational resources to the early stages is beneficial for enabling the network to better capture the geometric information carried by the raw point clouds. Based on this observation, we set the stage compute ratio of the four stages to $[3,1,1,1]$. It is worth noting that with this depth blocks setting, our modality-aware transformer backbone network achieved promising performance improvements in terms of 3.2 point in Success and 2.5 point in Precision metric, while significantly reducing the computational cost by half the GFLOPs compared to the baseline method (from 0.37 to 0.18).
Differences with Previous Study. Our results differ significantly from previous studies [44, 14], which were focused on LiDAR-based object detection tasks using CNN architecture with local receptive fields. In contrast, we focus on 3D SOT tasks within local search area and design a transformer architecture with global receptive fields. Therefore, by allocating more computational resources to the 1st stage, our network is able to capture the enough global-aware geometric information of the input point cloud. This is different with CNN-based backbone, which require more computational resources allocated to both the 1st and 2nd stages. As a result, our modality-aware transformer-based backbone design proves to be more effective.
IV-C Activation Function Selection for Point Cloud
[…]
$ReLU(x)=\max(0,x)$ (2)
$LeakyReLU(x)=\begin{cases}x&\text{if }x>0,\\ \alpha x&\text{if }x\leq 0.\end{cases}$ (3)
There is a discrepancy between the choice of activation functions natural language processing (NLP) and vision architectures. Rectified Linear Unit (ReLU)[48] is extensively used in convolutional networks (ConvNets) due to its simplicity and efficiency. With the Transformer[26] merging, Gaussian Error Linear Unit (GELU)[49] is considered as a smoother variant of ReLU and has been employed in state-of-the-art Transformer models, such as Google’s BERT [50], OpenAI’s GPT-3 [51], and Vision Transformers [35, 34], demonstrating superior performance across various language and vision tasks. However, unlike image or text data, point cloud input data contains negative values that represent corresponding geometric property. We assume that directly applying ReLU or GELU activation functions to point clouds would suppress the negative values and result in information loss. To address this issue, we explored the Leaky Rectified Linear Unit (LeakyReLU)[52], which can address the issue of neuron death and preserve more information within the range of values less than zero. As shwon in Eq. 1,2,3 The GELU and ReLU functions exhibit saturation behavior when the input values are less than zero, effectively reducing the information content. In contrast, the LeakyReLU function preserves the negative range of input values, enabling the retention of valuable information.
TABLE III: Influence of different activation functions on Car category.Bold denotes the best result
PFE Method Success Precision
GELU (SMAT[16]44footnotemark: 4) 68.9 79.6
ReLU 69.9 80.7
LeakyReLU 71.2 (+2.3) 81.8 (+2.2)
Figure 5: The impact of different activation functions on the reduction of the Epoch Loss.Inparticular, the LeakyReLU activation function has shown a more pronounced decrease in loss compared to the other two functions under consideration.
Therefore, we simply replace the GeLU activatation function in SMAT with LeakyReLU function. As shown in Tab III and Figure 5, using LeakyReLU function has better performance (+2.3 in Success and 2.2 in Precision) and lower loss value than other functions, highlighting the significant impact of activation function choice on 3D single object tracking tasks based on point cloud modality.
TABLE IV: Overall performance comparison on the KITTI dataset. The best two results are highlighted in red, blue.
Method Publication Modality Paradigm View Mean-by class Mean-by frame
Success Precision Success Precision
$M^{2}$-Track  [21] CVPR 2022 LIDAR Motion Point 61.9 80.2 60.6 80.1
SC3D[10] CVPR 2019 LIDAR Similarity Point 35.4 53.3 31.2 48.5
P2B[2] CVPR 2020 Point 39.5 53.9 42.4 60.0
3D-SiamRPN[1] IEEE Sensors J Point 43.8 58.6 46.6 64.9
MLVSNet[53] ICCV2021 Point 44.1 60.3 45.7 66.7
BAT[4] ICCV 2021 Point 47.2 65.1 51.2 72.8
PTT[3] IROS 2021 Point 48.4 63.4 55.1 74.2
LTTR[5] BMVC 2021 Voxel 50.0 67.4 48.7 65.8
V2B[22] NeurIPS 2021 Point+Voxel 52.4 65.6 58.4 75.2
PTTR[6] CVPR 2022 Point 58.4 77.8 57.9 78.1
CMT[27] ECCV 2022 Point 57.2 76.0 59.4 77.6
STNet[7] ECCV 2022 Point 63.4 81.4 61.3 80.1
C2FT[54] RAL 2022 Point 51.6 67.7 57.2 76.4
MLSET[55] RAL 2023 Point 54.2 68.9 59.6 78.4
SMAT[16] RAL 2022 Pillar 56.7 76.1 60.4 79.5
PillarTrack Ours Pillar 63.7 +7.0 78.9 +2.8 65.3 +4.9 82.2 +2.7
TABLE V: Detailed per-class performance on the KITTI dataset. The best two results are highlighted in red, blue.
Method Car-6424 Pedestrian-6088 Van-1248 Cyclist-308
Success Precision Success Precision Success Precision Success Precision
$M^{2}$-Track [21] 64.8 79.2 57.6 83.6 50.4 64.8 74.9 93.3
SC3D [10] 41.3 57.9 18.2 37.8 40.4 47.0 41.5 70.4
P2B [2] 56.2 72.8 28.7 49.6 40.8 48.4 32.1 44.7
3D-SiamRPN[1] 58.2 76.2 35.2 56.2 45.7 52.9 36.2 49.0
MLVSNet[53] 56.0 74.0 34.1 61.1 52.0 61.4 34.3 44.5
BAT [4] 60.5 77.7 42.1 70.1 52.4 67.0 33.7 45.4
PTT [3] 67.8 81.8 44.9 72.0 43.6 52.5 37.2 47.3
LTTR [5] 65.0 77.1 33.2 56.8 35.8 45.6 66.2 89.9
V2B [22] 70.5 81.3 48.3 73.5 50.1 58.0 40.8 49.7
PTTR [6] 65.2 77.4 50.9 81.6 52.5 61.8 65.1 90.5
CMT[27] 70.5 81.9 49.1 75.5 54.1 64.1 55.1 82.4
STNet [7] 72.1 84.0 49.9 77.2 58.0 70.6 73.5 93.7
C2FT [54] 67.0 80.4 48.6 75.6 53.4 66.1 38.0 48.7
MLSET [55] 69.7 81.0 50.7 80.0 55.2 64.8 41.0 49.7
SMAT [16] 71.9 82.4 52.1 81.5 41.4 53.2 61.2 87.3
PillarTrack (Ours) 74.2 +2.3 85.1 +2.7 59.7+7.6 84.7 +3.2 43.4 +2.0 51.7 77.4 +16.2 94.2 +6.9
V Experiments
KITTI Dataset. We conduct training and evaluating for our task using the KITTI [56] datasets. We follow the same setting in baseline model[16] to divide the sequences into training, validation, and testing splits, where 0-16 for training, 17-18 for validation and 19-20 for testing.
NuScenes Dataset. The NuScenes[57] dataset comprises 1000 scenes, which are allocated into 700/150/150 scenes for training, validation, and testing, respectively. Each scene is populated with approximately 300,000 points and is accompanied by comprehensive 360-degree viewpoint annotations. In our research, we adhere to the dataset configuration established by Open3DSOT[21] for the training and evaluation of our methodology. We also conduct a comparative analysis with the baseline model[16]to evaluate the efficacy and superiority of our designs.
Evaluation metrics. We use One-Pass Evaluation (OPE) [58] to measure Success and Precision. The Success measure the 3D IoU between the predicted box and the ground truth box, the Precision measures the AUC of the distance error between the center of the predicted and ground truth box.
Implementation details. We follow our baseline model[16] and employ PVTv2-b2 [37] as backbone, adhering to its original configuration. The specific settings include a head number of [1, 2, 5, 8], a depth configuration of [3, 1, 1, 1], the expansion ratio for the feed-forward layer set to [8, 8, 4, 8] and a channel number configuration of [64, 128, 320, 512]. In KITTI dataset, we set [0.1, 0.1, 4] meters as pillar size and a search area of [-3.2, -3.2, -3, 3.2, 3.2, 1] meters for the car category. In nuScenes dataset, we set [0.075, 0.075, 8] meters as pillar size and a search area of [-3.2, -3.2, -5, 3.2, 3.2, 3] meters for the car category. Additionally, the LeakyReLU activation is employed with a negative slope value of 0.01.
Training details. On the KITTI dataset, we train PillarTrack with AdamW[59] optimizer with an initial learning rate of 0.0001, a weight decay of 0.05, and a batch size of 16 on a NVIDIA 3090 GPU. The total number of training epochs was set to 72. For the NuScenes dataset, the training was conducted over 18 epochs, with a learning rate decay by a factor of 10 at epoch 14.
Method Paradigm Car-64159 Pedestrian-33227 Truck-13587 Trailer-3352 Bus-2953 Mean by class Mean by frame
Success Precision Success Precision Success Precision Success Precision Success Precision Success Precision Success Precision
$M^{2}$-Tracker [21] Motion 55.85 65.09 32.10 60.92 57.36 59.54 57.61 58.26 51.39 51.44 50.86 59.05 49.23 62.73
SC3D [10] Similarity 22.31 21.93 11.29 12.65 30.67 27.73 35.28 28.12 29.35 24.08 25.78 22.90 20.70 20.20
P2B [2] 38.81 43.18 28.39 52.24 42.95 41.59 48.96 40.05 32.95 27.41 38.41 40.90 36.48 45.08
PTT [12] 41.22 45.26 19.33 32.03 50.23 48.56 51.70 46.50 39.40 36.70 40.38 41.81 36.33 41.72
BAT [4] 40.73 43.29 28.83 53.32 45.34 42.58 52.59 44.89 35.44 28.01 40.59 42.42 38.10 45.71
GLT-T [60] 48.52 54.29 31.74 56.49 52.74 51.43 57.60 52.01 44.55 40.69 47.03 50.98 44.42 54.33
PTTR [6] 51.89 58.61 29.90 45.09 45.30 44.74 45.87 38.36 43.14 37.74 43.22 44.91 44.50 52.07
SMAT [16] 43.51 49.04 32.27 60.28 44.78 44.69 37.45 34.10 39.42 34.32 39.49 44.49 40.20 50.92
PillarTrack 47.12 57.72 34.18 64.93 54.82 54.41 57.70 54.63 44.68 40.73 47.70 54.48 44.59 58.86
Improvement +3.61 +8.68 +1.91 +4.65 +10.04 +9.72 +17.25 +20.53 +5.26 +6.41 +8.21 +9.99 +4.39 +7.94
TABLE VI: Performance comparison on the NuScenes dataset. The best two results are highlighted in red, blue.
V-A Quantitative results on KITTI.
As shown in Table IV and V, the proposed PillarTrack demonstrates significant performance improvements compared to the baseline method (SMAT [16]). Specifically, PillarTrack achieved a performance improvement of 7.0% 2.8% and 4.9% 2.7% in Success and Precision metrics in terms of mean by-class and mean-by frame setting, respectively.
In the categories of the Car, Pedestrian, Van, and Cyclist, PillarTrack outperformed SMAT by 2.3%, 7.6%, 2.0%, and 16.2% in Success, respectively. These improvements can be attributed to the effective designs of PillarTrack. Firstly, PE-PFE module could stabilize network training without information loss. Secondly, the modality-aware transformer-basde backbone focuses on capturing more geometric information from point clouds and enhance the feature representation capability of pillars. Lastly, LeakyReLU activation retains more information and help network optimization. Besides, PillarTrack also surpassed the motion-based method MMTrack [21] in the categories of Car, Pedestrian, and Cyclist, achieving state-of-the-art performance on the KITTI dataset. However, for Van, our method only has 43.4% in Success. We believe there are two main reasons for this. Firstly, the point clouds of vans are concentrated on the object surfaces, and after pillar encoding, they lack distinctive geometric features, making them more challenging to track accurately. Secondly, the limited number of training samples for the Van category may have constrained the model’s performance, as it might not have captured the full range of variations and characteristics exhibited by vans in the real world.
V-B Quantitative results on NuScenes.
As illustrated in Table VI, our PillarTrack achieves the second-best performance across the five categories on the nuScenes dataset. Compared to previous methods, M2-Tracker[21] adopts a different motion-based paradigm for object tracking, demonstrating superior results compared to the similarity-based paradigm. Meanwhile, our PillarTrack attains the best results among the similarity-based methods. Furthermore, compared to the baseline method[16], PillarTrack has demonstrated significant improvements in each category. Specifically, PillarTrack has enhanced the mean class-wise and frame-wise Success and Precision metrics by 8.21% / 9.99% and 4.39% / 7.94%, respectively. In terms of Success for each category, PillarTrack achieves notable improvements of 3.61% for Car, 1.91% for Pedestrian, 10.04% for Truck, 17.25% for Trailer, and 5.26% for Bus. Importantly, the Truck and Trailer categories demonstrate significantly enhanced performance compared to baseline. However, such improvements were not observed in the Van category in the KITTI dataset. We attribute this to two reasons: (1) The KITTI dataset only provides front-view annotations. Under the front-view perspective, the model is constrained to learn feature representations specific to the truck from a single viewpoint, which limits the acquisition of comprehensive geometric information about the Van. As a result, the Siamese network based on appearance matching struggles to capture the key geometric information of the Van. Conversely, the NuScenes dataset provides 360-degree surround view point cloud data with corresponding annotations. This enables superior learning of both appearance and geometric representations, leading to significant performance enhancement by effectively utilizing multi-view point cloud data. (2) The KITTI dataset contains a limited number of annotated van samples (1248), whereas the NuScenes dataset offers a larger sample size (13587/3352). Within our effective framework design, this facilitates richer feature representation learning, leading to performance improvements. Besides, our performance is slightly lower than that of PTTR on Car category. We think this is because PTTR utilizes both point-based and pillar-based representations to encode the input point cloud, resulting in superior performance, particularly in terms of the Success metric.
V-C Visualization results.
Figure 6: Visualization results comparison of our PillarTrack and baseline method (SMAT[16]) on the categories of Car, Pedestrian and Cyclist in KITTI dataset.
To comprehensively evaluate the performance of our method, particularly its capability to handle point clouds of varying sparsity, we conduct visualization on the different category, encompassing both sparse and dense scenes. Through comparative experiments on the initial frame and subsequent sequences, our method demonstrated superior tracking accuracy in both sparse and dense point cloud environments compared to the baseline SMAT method, highlighting ours enhanced adaptability and robustness. In the Cyclist category, our method was able to consistently track the target, whereas SMAT occasionally lost the target. For the Pedestrian category, our method maintained stable tracking from the outset, while SMAT failed to accurately identify the target initially. The visualization results in Figure 5 further substantiate the effectiveness and superior performance of our method across various scenarios.
V-D Running Speed
We calculated the average running time of all test frames for car to measure PillarTrack’s speed. PillarTrack model achieved 27 FPS on a single NVIDIA 3080 GPU, while baseline model SMAT has a 18.2 FPS in the inference phase. We did not introduce extra computational parameters, and we further reduced the computational overhead on the backbone while increasing the model’s performance. This once again demonstrates the effectiveness of our PillarTrack. Furthermore, our proposed PillarTrack can leverage quantization techniques[61, 62] to accelerate the runtime. We leave this problem for future research.
VI CONCLUSIONS
In this work, we proposed PillarTrack, a pillar-based 3D single-object tracking framework that decreases computation overhead while improving performance. The framework introduces a Pyramid-type Encoding Pillar Feature Encoder (PE-PFE) and a modality-aware Transformer-based backbone, and a simply replacement of activation function. Our extensive experiments on the KITTI dataset demonstrate PillarTrack’s superior performance, achieving an better trade-off between speed and accuracy. Based on above designs and its efficiency, We hope to inspire the community to rethink the design of network for 3D single object tracking based on point clouds.

### v2

@@
-PillarTrack: Redesigning Pillar-based Transformer Network for Single Object Tracking on Point Clouds
+PillarTrack: Boosting Pillar Representation for Transformer-based 3D Single Object Tracking on Point Clouds
 Abstract
 LiDAR-based 3D single object tracking (3D SOT) is a critical issue in robotics and autonomous driving.
-It aims to obtain accurate 3D BBox from the search area based on similarity or motion.
-However, existing 3D SOT methods usually follow the point-based pipeline, where the sampling operation inevitably leads to redundant or lost information, resulting in unexpected performance.
-To address these issues, we propose PillarTrack, a pillar-based 3D single object tracking framework.
-Firstly, we transform sparse point clouds into dense pillars to preserve the local and global geometrics.
-Secondly, we introduce a Pyramid-type Encoding Pillar Feature Encoder (PE-PFE) design to help the feature representation of each pillar.
-Thirdly, we present an efficient Transformer-based backbone from the perspective of modality differences.
-Finally, we construct our PillarTrack tracker based above designs.
-Extensive experiments on the KITTI and nuScenes dataset demonstrate the superiority of our proposed method.
-Notably, our method achieves state-of-the-art performance on the KITTI and nuScenes dataset and enables real-time tracking speed.
-We hope our work could encourage the community to rethink existing 3D SOT tracker designs.
+Existing 3D SOT methods typically adhere to a point-based processing pipeline, wherein the re-sampling operation invariably leads to either redundant or missing information, thereby impacting performance.
+To address these issues, we propose PillarTrack, a novel pillar-based 3D SOT framework.
+First, we transform sparse point clouds into dense pillars to preserve the local and global geometrics.
+Second, we propose a Pyramid-Encoded Pillar Feature Encoder (PE-PFE) design to enhance the robustness of pillar feature for translation/rotation/scale.
+Third, we present an efficient Transformer-based backbone from the perspective of modality differences.
+Finally, we construct our PillarTrack based above designs.
+Extensive experiments show that our method achieves comparable performance on the KITTI and NuScenes datasets, significantly enhancing the performance of the baseline.
 I INTRODUCTION
-3D Single object tracking (3D SOT) has a diverse applications in autonomous driving and robotics.
-Given the initial state (appearance and location) of specific target in the first frame, 3D single object tracking aims to estimate its 3D state across subsequent frames.
-Existing LiDAR-based 3D single object tracking methods [1, 2, 3, 4, 5, 6, 7] generally follow the Siamese paradigm from 2D visual object tracking, aiming to achieve a trade-off between runtime and accuracy.
-Given the sparse and irregular input point cloud, these methods initially leverage PointNet families [8, 9] to learn point-wise discriminative representation, and then obtain point-wise similarity with a feature aggregation module.
+3D single object tracking (3D SOT) has a wide applications in robotics, such as a robot needs to track their master to complete material handling tasks in factories.
+Given the initial state (size and location) of target in the first frame, 3D single object tracking aims to estimate its 3D state across subsequent frames.
+Existing LiDAR-based 3D single object tracking methods [1, 2, 3, 4, 5, 6, 7] generally follow the Siamese paradigm from 2D visual object tracking [8, 9], aiming to achieve a trade-off between runtime and accuracy.
+Given the sparse and irregular input point cloud, these methods initially leverage PointNet families [10, 11] to learn point-wise representation, and then obtain point-wise similarity with a similarity module.
 Finally, they estimate the state of specific targets based on these similarity features.
 SC3D [12] stands as the pioneering LiDAR-based 3D Siamese tracker, leveraging an efficient auto-encoder based on PointNet[10] for point-wise feature encoding.
-Subsequently, approaches such as P2B [2], 3D-SiamRPN[1], BAT [4] and PTT [3, 12] employ PointNet++ [9] to extract more effective point-wise representation and achieve superior tracking performance.
-However, one common issue with above point-based 3D SOT methods is the need for re-sampling the input points to a fixed number.
+Subsequently, approaches such as P2B [2], 3D-SiamRPN[1], BAT [4] and PTT [3, 13] employ PointNet++ [11] to extract point-wise representation and achieve superior tracking performance.
+However, one common issue with above point-based 3D SOT methods is the requirement for re-sampling the input points to a fixed number.
 For instance.
-P2B [2] requires the input points to be re-sample to 1024 for the search area and 512 for the template point cloud in order to align the network input.
+P2B [2] re-samples the input points from 1024 to 512.
 This re-sample operation inevitably introduces the potential for redundant or lost information, which can adversely affect performance.
-Additionally, the reliance on point query/retrieval in 3D space (e.g., PointNet++ [9]) for the re-sampling process may not be friendly to efficient hardware implementation.
-Figure 1: Comparison with other 3D SOT methods on KITTI dataset. We categorize methods based on the Backbone architecture and report their performance on Success and Precision.
-To this end, inspired by the recent advancements in pillar-based 3D object detectors [13, 14, 15] due to their real-time speed and high performance.
+Figure 1: Comparison with other 3D SOT methods on KITTI dataset. We classify methods based on the backbone architecture and report performance on Success and Precision.
+To this end, inspired by the recent advancements in pillar-based 3D object detectors [14, 15] due to their real-time speed and high performance.
 Therefore, we convert the sparse and irregular point clouds into dense and regular pillar representations.
-It is noteworthy that the 3D SOT task can be regarded as a specific 3D object detection task within a local search area, with the additional template point prior information.
-Therefore,it is natural to leverage advanced principles from 3D point cloud detection to improve point cloud tracking tasks.
 Specifically, pillar-based point cloud representation possesses following advantages: (1) Pillar representation is dense and ordered, facilitating seamless integration with advanced 2D image-based techniques without much modification.
@@
 Particularly, pillar representation naturally aligns with the high demanding real-time requirements of 3D trackers, making it well-suited for tracking tasks.
-Currently, few works [16, 17] are dedicated to pillar-based 3D SOT with point clouds.
+Currently, there are few works [16, 17] are dedicated to pillar-based 3D SOT tasks.
 SMAT [16] addresses this gap by transforming sparse 3D point clouds into dense pillars and then utilizing a transformer-based encoder for multi-scale feature extraction and interaction, yielding promising results.
-However, SMAT directly adopts a simplistic pillar encoding module from PointPillars [18] limits the quality of the pillar representations.
-Besides, their backbone directly follows the design principle of vision transformers tailored for RGB image, which may be sub-optimal for point cloud modalities.
-PTTR++ [17] introduces a fusion of point-wise and pillar-wise views to further enhance the performance of PTTR [6].
-Nevertheless, PTTR++ still involves a re-sampling operation within its pipeline, which presents paratical deployment challenges.
-In this paper, we propose PillarTrack, a pillar-based 3D single-object tracking framework, aiming to improve tracking performance while enhancing the inference speed.
-Firstly, we transform sparse and unordered point clouds into dense and regular pillar representations to reduce information loss caused by re-sampling operation and propose a Pyramid-type Encoding Pillar Feature Encoder (PE-PFE) design to help the feature learning of each pillar.
-Secondly, we argue the design of existing Transformer-based backbone from the perspective of modality differences and propose an modality-aware Transformer-based backbone specifically tailored for point cloud modalities, achieving higher performance with fewer GFLOPs.
-Finally, we construct our PillarTrack network.
-Extensive experiments on the KITTI and nuScenes dataset demonstrate the superiority of our proposed method.
-As shown in Figure  1, our approach achieves state-of-the-art performance on the KITTI and nuScenes dataset, striking a better balance between speed and accuracy, and catering to diverse practical needs.
+However, SMAT utilizes a naive pillar encoding module derived from PointPillars [18], which retains the substantial numerical disparities between point cloud channels.
+This approach doesn’t fully leverage the intrinsic structure of the point cloud data and limit the representation robustness of pillar.
+Furthermore, its backbone directly adheres to the design principles of visual transformers [19, 20] that are tailored for RGB images, which is not optimal for the point cloud modality.
+PTTR++ [17] introduces point-wise and pillar-wise view fusion to further enhance the performance of PTTR [6].
+Nonetheless, PTTR++ still incorporates a re-sampling operation in its process, introducing additional computational overhead and posing challenges for practical deployment.
+In this paper, we propose PillarTrack, a pillar-based 3D single-object tracking framework, with the objective of augmenting tracking performance and accelerating inference speed, thereby facilitating deployment.
+Firstly, to reduce information loss caused by re-sampling operation, we propose a Pyramid Encoding Pillar Feature Encoder (PE-PFE) design to transform sparse point clouds into dense and robust pillar representations.
+Secondly, we argue the design of existing Transformer-based backbone from the perspective of modality differences and propose an modality-aware Transformer-based backbone specifically tailored for point cloud modalities, achieving higher performance with fewer computation resource.
+Finally, we construct our PillarTrack based on above designs.
+Extensive experiments on the KITTI and nuScenes dataset demonstrate the superiority of our method.
+As shown in Figure 1, PillarTrack achieves comparable performance on the KITTI and nuScenes dataset, striking a better balance between speed and accuracy, and catering to diverse practical needs.
 In conclusion, the main contributions of this paper are as follows:
-$\bullet$ PE-PFE: A Pyramid-type Encoding Pillar Feature Encoder (PE-PFE) design to encode the point coordinate of each pillar with pyramid-type representation, and bring the performance gain without extra computational overhead.
-$\bullet$ Modality-aware Transformer-based Backbone: A more suitable for point cloud modalities, aiming to enhance feature representation.
-This design involves a simple adjustment of computational resources allocated to the front-end of the backbone, allowing for capturing more semantic details of the input point cloud.
-$\bullet$ SOTA and Open-source: Experiments on KITTI and nuScenes dataset show our method achieve the state-of-the-art performance.
-Besides, we will open source our code to the research community in https://github.com/StiphyJay/PillarTrack.
+$\bullet$ PE-PFE: A Pyramid-Encoded Pillar Feature Encoder (PE-PFE) is designed to encode the point coordinate of each pillar with pyramid-type representation.
+This approach not only enhances performance without incurring additional computational overhead but also improves robustness against rotation, translation, and scale variations.
+$\bullet$ Modality-aware Transformer-based Backbone: A tailored for point cloud modalities backbone design, aiming to enhance pillar representation.
+It involves an adjustment of computational resources allocated to the front-end of the backbone, enabling the capture of more inherent semantic details from the input point cloud.
+$\bullet$ Comparable and Open-source: Extensive experiments on KITTI and nuScenes dataset shown the comparable performance, significantly improving upon the baseline approaches.
+Besides, we open source our code to the community in https://github.com/StiphyJay/PillarTrack.
 II Related Work
-II-A Single Object Tracking on Point Clouds
-Benefiting from the nature of LiDAR that are less sensitive to illumination changes and capture accurate distance information, numerous works[10, 19, 20, 1, 2, 4, 21, 22] in LiDAR-based 3D Single Object Tracking (3D SOT) have merged.
-SC3D[10] leverages the Kalman filter for proposals generation and select the most suitable candidate based on Siamese network[23].
-3D-SiamRPN [1] combines 3D Siamese network with 3D RPN network for object tracking.
-P2B[2] introduces target-specific features and utilizes VoteNet[24] to estimate target center.
-Based on the P2B pipeline, BAT[4] incorporated Box-aware information to enhance similarity feature.
-Besides, V2B[22] introduces Voxel-to-BEV network for pixel-wise regression of 3D targets, M2-Track [21, 25] adopts a motion-centric paradigm to address the issue of 3D SOT and achieves remarkable performance.
-Inspired by the success of transformers[26], several methods [3, 12, 5, 6, 27, 7, 28, 29] have merged.
+Single Object Tracking on Point Clouds.
+Leveraging the advantages of LiDAR, which is less sensitive to illumination changes and captures accurate distance information, numerous works[12, 21, 1, 2, 4, 22, 23] in LiDAR-based 3D Single Object Tracking (3D SOT) have merged.
+SC3D[12] utilizes a Kalman filter for proposals generation and selects the most suitable candidate via a Siamese network.
+3D-SiamRPN [1] combines 3D Siamese network with 3D RPN network for tracking.
+P2B[2] introduces target-specific features and employs VoteNet[24] to estimate the target center.
+Based on this, BAT[4] incorporates box-aware information to enhance similarity feature.
+M2-Track [22, 25] adopts a motion-centric paradigm that yields impressive results.
+Inspired by the success of transformers[26], several transformer-based methods [3, 13, 5, 6, 27, 7, 28, 29] have merged.
 PTT [3, 13] proposed Point-Track-Transformer module to assign weights to crucial point cloud features.
-LTTR[5] proposes a Transformer-based feature fusion module to learn semantic context among different region.
 PTTR[6] employs an attention-based feature matching mechanism to integrate target clues in similarity.
 STNet[7] develops an iterative coarse-to-fine correlation network for robust correlation learning.
-CXTrack[29] adopts transformer to extract richer contextual information from consecutive frames, enhancing localization accuracy.
-Recently, SMAT[16] encodes global similarity based on attention in the BEV to alleviate the impact of point sparsity.
-However, their backbone still partially inherits the design principles from ResNet[30] tailored for the image modality, which may be sub-optimal for point clouds modality.
-Figure 2: The architecture of our PillarTrack network.Given the template and search area, we first use PE-Pillar to extract multi-scale features respectively.
-Then, MAE computes similarity by attention at each feature scale and then fuses the multi-scale similarity feature.Finally, we apply the detection head on the feature fusion map to localize the target.
-II-B Pillar-based Object Detection on Point Clouds
-Grid-based point cloud 3D detectors[18, 31, 13, 14, 15] generally transform unordered points into regular pillar/voxel and then utilize 2D/3D CNN to learn representation.
-VoxelNet [32] stands as a pioneering work that leverages Voxel Feature Extractor (VFE) and 3D CNN for 3D object detection.
-Subsequently, SECOND [33] employs 3D sparse convolution on non-empty voxels to accelerate training and inference.
-Different from above approaches, PointPillars [18] innovatively converts point clouds into pillars and leverages 2D CNN to enhance detection speed.
-After that, CenterPoint[31] proposes a center-based head design, achieving milestone performance.
-Furthermore, PillarNet[13] improves the accuracy of pillar-based 3D detectors through an ”encoder-neck-head” architecture, achieving superior performance while maintaining real-time speed.
-Based on PillarNet[13], FastPillars[14] proposes a backbone design suitable for point cloud modality and considers deployment-friendly ability, resulting in new state-of-the-art.
-The success of these pillar-based 3D detectors greatly motivates us to explore the potential of simple pillar-based representation in 3D SOT tasks.
-II-C Vision Transformer
-Transformer[26] is first proposed in natural language processing (NLP).
-Subsequently, owing to its remarkable ability for global modeling, the Transformer has shown significant success in image recognition tasks.
-ViT[34] leverage the Transformer structure to classify images by splitting them into patches.
-SwinTransformer[35] proposes a shifted windows attention mechanism, achieving excellent performance across various image tasks.
-PVT[36] and PVTv2[37] utilize a progressive shrinking pyramid to construct a visual Transformer backbone.
-In the 3D Vison, Point Transformer[38] introduces a specialized Point Transformer layer tailored for processing unordered point clouds.
-PCT[39] utilizes 3D coordinates of points to generate position embedding and employs an attention module to extract point features from local neighborhoods.
-These efforts showcase outstanding performance in point cloud shape classification or part segmentation.
-For LiDAR-based 3D detection tasks, recent approaches[40, 41] first transform point clouds into regular pillars/voxels and then utilize dynamic sparse windows on BEV feature for feature extraction, yielding remarkable performance.
-In contrast, we employ the Transformer for global modeling on the pillarized feature, aiming to explore the potential of the Transformer in the object tracking task from a BEV perspective.
+GLT-T[28] proposes a global-local Transformer voting scheme to generate higher quality 3D proposals.
+Recently, SMAT[16] encodes global similarity based on attention to alleviate the point sparsity.
+However, their backbone still partially inherits the design principles from ResNet[30] tailored for the image modality, which is sub-optimal for point clouds.
+Therefore, we argue their design and propose a modality-aware transformer-based backbone tailored for point cloud modality.
+Figure 2: The architecture of our PillarTrack network.
+Given the template and search area, we first use PE-PFE to extract multi-scale features respectively.
+Then, the backbone extract features at different feature scale and fuses the multi-scale similarity feature.
+Finally, we apply the detection head on the feature fusion map to localize the target.
+Vision Transformer.
+The Transformer architecture [26], originally proposed for natural language processing (NLP), has demonstrated significant success in image recognition due to its global modeling capabilities.
+The Vision Transformer (ViT)[31] classifies images by dividing them into patches, while Swin Transformer[32] introduces a shifted window attention mechanism, achieving excellent performance across various visual tasks.
+PVT [19] and PVTv2 [20] employ a progressive shrinking pyramid to construct a visual Transformer backbone.
+In 3D vision, Point Transformer [33] features a specialized Point Transformer layer for processing unordered point clouds.
+These methods have shown outstanding performance in point cloud shape classification and part segmentation.
+In contrast, our approach explores the potential of the Transformer’s global modeling on pillar features in the object tracking task.
 III Preliminary
-III-A Problem Statement
-In the task of the 3D Single Object Tracking (3D SOT), we initially define target point cloud (i.e., template) $\mathbf{P_{t}}=\{\mathbf{p}_{ti}=[x_{i},y_{i},z_{i},r_{i}]^{T}\in\mathbb{R}^{N\times 4}\}$, where $x_{i},y_{i},z_{i}$ denote the coordinate values of each point along the axes X, Y, Z, respectively, and $r_{i}$ represents the laser reflection intensity.
+Motivation.
+Our research endeavors to investigate a real-time, high-performance, pillar-based approach for 3D single-object tracking.
+To achieve this, we leverage the compact pillar representation and design a strong Transformer-based backbone that enhances pillar feature representation with reduced computational resources from the perspective of inherent modality differences.
+Regarding the feature integration and similarity computation module, we follow the original design of baseline[16], which has shown remarkable efficacy.
+We hope that our designs can offer a novel perspective in 3D SOT, encouraging the community to revisit existing transformer-based methods.
+Task Definition.
+In the task of the 3D Single Object Tracking (3D SOT), we initially define target point cloud (i.e., template) $\mathbf{P_{t}}=\{\mathbf{p}_{i}^{t}=[x_{i},y_{i},z_{i},r_{i}]^{T}\in\mathbb{R}^{N\times 4}\}$, where $x_{i},y_{i},z_{i}$ denote the coordinate values of each point along the axes X, Y, Z, respectively, and $r_{i}$ represents the laser reflection intensity.
 Our goal is to locate the the 3D bounding box (3D BBox) of this target in the search area $\mathbf{P_{t}}=\{\mathbf{p}_{si}=[x_{j},y_{j},z_{j},r_{j}]^{T}\in\mathbb{R}^{M\times 4}\}$ frame by frame.
@@
 By applying the displacement and heading angle from the previous frame’s 3D BBox, we can locate the target in the current frame.
-III-B Baseline Architecture Overview
-As SMAT represents the first attempt at a pillar-based 3D Siamese tracker and has shown promising performance, we choose SMAT as our baseline and provide a brief overview of its structure.
+Baseline Overview.
+As SMAT is the first attempt at a pillar-based 3D Siamese tracker and has shown promising results, we choose SMAT as our baseline and briefly review its structure.
 The core components of SMAT include: 1) a naive pillar feature encoding module derived from Pointpillars, which transforms raw point cloud data into a pseudo-image; 2) a Transformer-based backbone network that computes similarity through a multi-head attention mechanism for the multi-scale features from both the template and search branches, capturing local and global information within the point cloud; 3) a neck module that integrates features from different scales and fuses multi-scale similarity features, enhancing the model’s perception of multi-scale information; 4) a decoder that predicts the position and orientation of the target object.
-These carefully designed features enable SMAT to effectively handle sparse point cloud data, offering an efficient solution for 3D Single Object Tracking.
-III-C Motivation
-Our research endeavors to investigate a real-time, high-performance, pillar-based approach for 3D single-object tracking.
-To achieve this, we leverage the compact pillar representation and design a strong Transformer-based backbone that enhances pillar feature representation with reduced computational resources from the perspective of modality differences.
-Regarding the feature integration and similarity computation module, we follow the original design of baseline method[16], which has shown remarkable efficacy.
-We hope that our design principles offer a novel perspective in 3D single object tracking, encouraging the community to revisit existing methodologies.
+These designs enable SMAT to effectively handle sparse point cloud, offering an efficient solution for 3D SOT.
 IV Method
-IV-A Pyramid-Encoded Pillar Feature Encoding Design
-Figure 3: The illustration of PE-PillarVFE design.
-We encode the channel information of the input point cloud in a pyramid-type fashion.
-This Pyramid-type encoding design allows the network to optimize more effectively without the input information loss.
+IV-A Pyramid-Encoded Pillar Feature Encoding
+Figure 3: The illustration of PE-PFE design.
+We encode the point coordinate of the input point cloud in a pyramid-like type.
+This Pyramid-like encoding design allows the network to optimize effectively without the input information loss.
+Figure 4: Illustration of the properties of PE-PFE.
+The red indicates the feature distribution after original PFE encoding, while the blue represents the feature distribution after PE-PFE encoding.
+(a) Original point cloud.
+(b) Point cloud with 1.2X scale.
+(c) Point cloud with 1.2m translation.
+(d) Point cloud with 45°$\degree$ rotation.
+Figure 5: The output features visualization of forth stage in t-SNE.(a) Visualization with the image blocks setting.
+(b) Visualization matching the total number of the original blocks.
+(c) Visualization after removing redundancies from the number of blocks.
 In pillar-based 3D detection methods[18, 14, 15], Pillar Feature Encoding (PFE) plays a crucial role, directly impacting the representation learning of BEV feature and final detection accuracy.
-Its significance extends to 3D single object tracking tasks as well.
-In SMAT, they utilize the PFE module from PointPillars [18] to extract pillar-wise feature.
+Its significance also extends to 3D single object tracking tasks.
+SMAT utilizes the PFE from PointPillars [18] to extract pillar-wise feature.
 Here, taking PFE module as an example, we illustrate the details of their input encoding.
-The points in each pillar are initially concatenated along the channel dimension to form $\hat{p_{i}}=\{[x_{i},y_{i},z_{i},r_{i},{x}_{i}^{m},{y}_{i}^{m},{z}_{i}^{m},{x}_{i}^{c},{y}_{i}^{c},{z}_{i}^{c}]\in{\mathbb{R}^{N_{v}\times 10}}\}$, where $[x_{i},y_{i},z_{i}]$ is the original point coordinates in the ego frame, $[{x}_{i}^{m},{y}_{i}^{m},{z}_{i}^{m}]$ denotes the offset of $p_{i}$ from the mean of the xyz coordinates of all points within the current pillar, and $[{x}_{i}^{c},{y}_{i}^{c},{z}_{i}^{c}]$ is the offset of $p_{i}$ from the current pillar center.
-However, there exist significant numerical differences in the point-wise feature.
-For example, the numerical range of the $[x_{i},y_{i},z_{i}]$ is $[-3.2,3.2]$ (i.e., the range of search area).
-However, the numerical range of $[{x}_{i}^{m},{y}_{i}^{m},{z}_{i}^{m}]$ and $[{x}_{i}^{c},{y}_{i}^{c},{z}_{i}^{c}]$ is $[-0.1,0.1]$, since they are relative offset within a pillar.
-It can be observed that there exist significant numerical differences (almost 30 $\times$) between the point-wise feature channels in PFE.
-Such huge disparities in input values pose significant challenges for network optimization  [42, 43].
-Therefore, we propose a Pyramid-type Encoding Pillar Feature Encoder (PE-PFE) design aimed at reducing the numerical differences between the input channels, enabling the network towards better optimization on numerically stable inputs.
-TABLE I: Tracking performance comparison of different PFE designs in the Car category in KITTI.
-PFE Method Success Precision
-PFE (SMAT [16] [^4] The model was trained on a setup comprising four NVIDIA GeForce RTX 3090 GPUs) 68.9 79.6
-PE-PFE(Ours) 70.8 (+1.9) 80.3 (+0.7)
-Figure 4: The loss comparison of different PFE designs.
-As shown in Figure 3, we encode the channel values of the input point cloud in a pyramid-type manner, and finally retain a residual value that cannot be evenly divided.
+The points in each pillar are initially concatenated along the channel dimension to form $\hat{p_{i}}=\{[x_{i},y_{i},z_{i},r_{i},{x}_{i}^{m},{y}_{i}^{m},{z}_{i}^{m},{x}_{i}^{c},{y}_{i}^{c},{z}_{i}^{c}]\in{\mathbb{R}^{N_{v}\times 10}}\}$, where $[x_{i},y_{i},z_{i}]$ is the original point coordinates in the ego frame, $N_{v}$ is the number of points within the pillar, $[{x}_{i}^{m},{y}_{i}^{m},{z}_{i}^{m}]$ denotes the offset of $p_{i}$ from the mean of the xyz coordinates of all points within the current pillar, and $[{x}_{i}^{c},{y}_{i}^{c},{z}_{i}^{c}]$ is the offset of $p_{i}$ from the current pillar center.
+However, there exist large numerical differences between the feature channels in PFE.
+For instance, the numerical range of $[{x}_{i}^{m},{y}_{i}^{m},{z}_{i}^{m}]$ and $[{x}_{i}^{c},{y}_{i}^{c},{z}_{i}^{c}]$ is $[-0.1,0.1]$, since they are relative offset within a pillar.
+Such huge disparities in input values pose significant challenges for network optimization [34] and can make model sensitive to disturbance, such as scaling, translation or rotation.
+To this end, we propose the Pyramid-Encoding Pillar Feature Encoder (PE-PFE), which is designed to minimize value discrepancies between channels, thereby facilitating improved optimization of numerically stable inputs.
+Besides, PE-PFE ensures that input point clouds maintain robustness against scale, translation, rotation.
+In detail, as shown in Figure 3, we encode the channel values of the input point cloud in a pyramid-type manner, and finally retain a residual value that cannot be evenly divided.
 The encoding values from each layer and the residual value are then cascaded together as point-wise features, which are then processed by the PFE module.
-It is important to note that this pyramid-type encoding design significantly reduces the numerical differences between the input data channels without sacrificing the information contained in the point cloud.
-This design helps the network achieve better optimization results.
-Furthermore, to demonstrate the effectiveness of our PE-PFE design compared to the naive PFE from PointPillars [18], we present the performance comparison in the Car category.
-As shown in Table I, our simple yet effective PE-PFE design yields 1.9 point gains in Success and 0.7 point improvement in Precision.
-As shown in Figure 4, the design of PE-PFE leads to lower loss values, enabling the network to find better parameters and achieve superior performance.
+Similar to the natural language processing (NLP) tokenization, the PE-PFE serves as a feature embedding that converts point cloud data into an information-dense representation.
+The embedding of PE-PFE is characterized by threes properties: scale-robustness, translation-robustness, and rotation-robustness.
+Scale robustness aids in the feature extraction of geometrical information for vehicles with identical shapes but varying sizes, while remaining insensitive to scale variations.
+Translation-robustness ensures that even with positional perturbations in the input point cloud, the distribution of feature encoding remains stable, facilitating reliable representation.
+Rotation-robustness assists in extracting point cloud with different orientation angle, enhancing the expression of semantics.
+Collectively, these properties contribute to the robustness and effectiveness of PE-PFE in feature extraction and geometric understanding within the 3D SOT tasks.
+For instance, as show in Figure 4, we apply different operations of 1.2X scale, 1.2m translation, and 45°$\degree$ rotation to the raw point cloud of a vehicle.
+The input feature distribution of original PFE exhibits significant discrepancies; however, after PE-PFE, the feature distributions remained more stable.
+We remain the exploration of PE-PFE on other Lidar-based tasks in the future work.
+For the effectiveness of PE-PFE, refer to Table V.
 IV-B Modality-aware Transformer-based Backbone
-Recent studies [44, 14, 45] have pointed out that the backbone designs commonly employed in image domain [46, 30, 35, 47] are sub-optimal for point cloud or depth images due to significant modality differences between RGB images and LiDAR point clouds.
-Particularly, backbone design in image domain tends to utilize more block layers in later stages to extract high-level semantic features for perception.
-For instance, ResNet-50[30] takes 6 residual blocks in the 3rd stage while the 1st stage uses 3 blocks for feature extraction.
-Similarly, Swin-Transformer [35] adopts 9 blocks in the 3rd stage while only 1 block is used in the 1st stage.
-However, we think that these design choices for distributing computations across stages were largely empirical and may not be well-suited for LiDAR point clouds.
+Recent studies [35, 15, 36] have pointed out that the backbone designs commonly employed in image domain [30, 32] are sub-optimal for point cloud or depth images.
+This is due to the significant modality differences between RGB images and LiDAR point clouds, where LiDAR inherently capture geometric semantics.
+Particularly, backbone design in image domain tends to utilize more computation burden in later stages to extract high-level semantic features for perception tasks.
+For instance, ResNet-50[30] takes 6 residual blocks in the 3rd stage while using 3 blocks in 1st stage for feature extraction.
+Similarly, Swin-T [32] adopts 9 blocks in the 3rd stage while only 1 block is used in the 1st stage.
+However, we argue that these design choices for distributing computations across stages were largely empirical and not be well-suited for LiDAR point clouds.
 Unlike RGB images, LiDAR can capture spatial distances and shapes of objects.
-This means that rich and accurate geometric information of objects is already explicitly encoded in point cloud.
-Therefore, we assume that instead of allocating excessive computational resources to model object geometries in later stages like ResNet/Swin-Transformer in RGB images, we propose reallocating more computational resources to the early stages to effectively incorporate the geometric information carried by raw point clouds.
-TABLE II: Impact of varying block quantities in the four stages of the backbone on the final results. Bold denotes the best performance.
+This means that rich and accurate geometric information of objects is already explicitly captured in raw point cloud coordinates.
+Therefore, we assume that instead of allocating excessive computational resources to model object geometries in later stages like ResNet/Swin-T in RGB images, we propose reallocating more computational resources to the early stages to effectively incorporate the geometric information carried inherenty by point clouds.
+To validate the aforementioned hypothesis for a transformer-based backbone, we performed t-SNE111t-SNE (t-distributed Stochastic Neighbor Embedding) is a machine learning algorithm designed for the visualization of high-dimensional data.
+It maps high-dimensional data to a lower-dimensional space while preserving relative distances and distribution structures, thereby facilitating exploratory data analysis and providing insights into the internal structure of the data.
+visualization [37] on the features processed at different stages.
+The effectiveness of feature extraction can be evaluated by observing cluster separation, shape, and compactness in the t-SNE visualization.
+Clearly separated, regular, and compact clusters suggest successful feature extraction.
+Initially, we visualized the features under the original blocks number setting using t-SNE, as indicated in the red box of Figure 5 (a).
+The visualization revealed significant overlap between clusters, irregular shapes, and lack of compactness, indicating that the features were not adequately processed under the original configuration.
+Therefore, keeping the total number of blocks constant, we allocated more computational resources to the earlier stages of the network.
+As shown in the red box of Figure 5 (b), this adjustment led to less overlap between clusters compared to the original setting, demonstrating some improvement and the learned diverse representation.
+However, the clusters still lacked compactness, leading us to hypothesize that there was redundant computation.
+To test this hypothesis, we conducted a comparative experiment with a ”3111” block configuration, as illustrated in the red box of Figure 5 (c).
+This setup resulted in clearly separated clusters, many of which exhibited an elliptical shape, suggesting that our hypothesis was preliminarily confirmed.
+In summary, the adjustments made to the allocation of computational resources across the stages of the transformer-based backbone have significantly improved the quality of feature extraction, as evidenced by the enhanced cluster separation and regularity observed in the t-SNE visualizations.
+TABLE I: Performance comparison on the KITTI dataset.
+L means Lidar-only methods.
+M and S are Motion-based and Similarity-based paradigms.
+The best two results are highlighted in red, blue.
+* indicates results produced based on the official implementation.
+Method Publication Paradigm View Car-6424 Pedestrian-6088 Van-1248 Cyclist-308 Mean-by class Mean-by frame
+Success Precision Success Precision Success Precision Success Precision Success Precision Success Precision
+$M^{2}$-Track*  [22] CVPR 2022 L+M Point 64.8 79.2 57.6 83.6 50.4 64.8 74.9 93.3 61.9 80.2 60.6 80.1
+SC3D[12] CVPR 2019 L+S Point 41.3 57.9 18.2 37.8 40.4 47.0 41.5 70.4 35.4 53.3 31.2 48.5
+P2B[2] CVPR 2020 Point 56.2 72.8 28.7 49.6 40.8 48.4 32.1 44.7 39.5 53.9 42.4 60.0
+3D-SiamRPN[1] IEEE Sensors J Point 58.2 76.2 35.2 56.2 45.7 52.9 36.2 49.0 43.8 58.6 46.6 64.9
+MLVSNet[38] ICCV2021 Point 56.0 74.0 34.1 61.1 52.0 61.4 34.3 44.5 44.1 60.3 45.7 66.7
+BAT[4] ICCV 2021 Point 60.5 77.7 42.1 70.1 52.4 67.0 33.7 45.4 47.2 65.1 51.2 72.8
+PTT[3] IROS 2021 Point 67.8 81.8 44.9 72.0 43.6 52.5 37.2 47.3 48.4 63.4 55.1 74.2
+LTTR[5] BMVC 2021 Voxel 65.0 77.1 33.2 56.8 35.8 45.6 66.2 89.9 50.0 67.4 48.7 65.8
+V2B[23] NeurIPS 2021 Point+Voxel 70.5 81.3 48.3 73.5 50.1 58.0 40.8 49.7 52.4 65.6 58.4 75.2
+PTTR[6] CVPR 2022 Point 65.2 77.4 50.9 81.6 52.5 61.8 65.1 90.5 58.4 77.8 57.9 78.1
+CMT[27] ECCV 2022 Point 70.5 81.9 49.1 75.5 54.1 64.1 55.1 82.4 57.2 76.0 59.4 77.6
+STNet[7] ECCV 2022 Point 72.1 84.0 49.9 77.2 58.0 70.6 73.5 93.7 63.4 81.4 61.3 80.1
+C2FT[39] RAL 2022 Point 67.0 80.4 48.6 75.6 53.4 66.1 38.0 48.7 51.6 67.7 57.2 76.4
+MLSET[40] RAL 2023 Point 69.7 81.0 50.7 80.0 55.2 64.8 41.0 49.7 54.2 68.9 59.6 78.4
+Baseline* [16] RAL 2022 Pillar 71.9 82.4 52.1 81.5 41.4 53.2 61.2 87.3 56.7 76.1 60.4 79.5
+PillarTrack* Ours Pillar 74.2 85.1 59.7 84.7 61.0 69.2 78.0 95.0 68.2 83.5 66.8 83.7
+Improvement +2.3 +2.7 +7.6 +3.2 +19.6 +16.0 +16.8 +7.7 +11.4 +7.2 +6.4 +4.2
+To further substantiate our hypothesis from the results, we systematically study computation allocation by adjusting the stage compute ratio in the transformer-based backbone of SMAT [16].
+It is worth noting that the setting of SMAT has the same depth number as ResNet-50, with a stage compute ratio of $[3,4,6,3]$.
+As shown in Table IV, we used the SMAT backbone as the baseline and varied the number of depth blocks from 0 to 4 with a stride of 1 in each stage.
+From the results, we can find that the performance is quite sensitive to the capacity of front stage, while being less sensitive to the later stages (2, 3, 4).
+This confirmed that allocating more computational resources to the early stages is beneficial for enabling the network to better capture the geometric information carried by the raw point clouds.
+Based on this observation, we set the stage compute ratio of the four stages to $[3,1,1,1]$.
+It is worth noting that with this depth blocks setting, our modality-aware transformer backbone network achieved promising performance improvements in terms of 3.2/in Success and 2.5/in Precision metric, while significantly reducing the computational cost by half the GFLOPs compared to the baseline method (from 0.37 to 0.18).
+It is noteworthy that an increase in computational load at the front end leads to a saturated performance.
+We surmise this is due to the saturation of the network’s parameters, where the excessive employment of parameters augments the learning burden on the network.
+This phenomenon naturally substantiates the existence of a performance bottleneck in the principle of computational resource allocation, illustrating that performance does not escalate linearly with additional computational resources.
+TABLE II: Performance comparison on the NuScenes dataset. The best two results are highlighted in red, blue.
+Method Paradigm Car-64159 Pedestrian-33227 Truck-13587 Trailer-3352 Bus-2953 Mean by class Mean by frame
+Success Precision Success Precision Success Precision Success Precision Success Precision Success Precision Success Precision
+$M^{2}$-Tracker [22] Motion 55.85 65.09 32.10 60.92 57.36 59.54 57.61 58.26 51.39 51.44 50.86 59.05 49.23 62.73
+SC3D [12] Similarity 22.31 21.93 11.29 12.65 30.67 27.73 35.28 28.12 29.35 24.08 25.78 22.90 20.70 20.20
+P2B [2] 38.81 43.18 28.39 52.24 42.95 41.59 48.96 40.05 32.95 27.41 38.41 40.90 36.48 45.08
+PTT [13] 41.22 45.26 19.33 32.03 50.23 48.56 51.70 46.50 39.40 36.70 40.38 41.81 36.33 41.72
+BAT [4] 40.73 43.29 28.83 53.32 45.34 42.58 52.59 44.89 35.44 28.01 40.59 42.42 38.10 45.71
+GLT-T [28] 48.52 54.29 31.74 56.49 52.74 51.43 57.60 52.01 44.55 40.69 47.03 50.98 44.42 54.33
+PTTR [6] 51.89 58.61 29.90 45.09 45.30 44.74 45.87 38.36 43.14 37.74 43.22 44.91 44.50 52.07
+Baseline [16] 43.51 49.04 32.27 60.28 44.78 44.69 37.45 34.10 39.42 34.32 39.49 44.49 40.20 50.92
+PillarTrack Ours 47.12 57.72 34.18 64.93 54.82 54.41 57.70 54.63 44.68 40.73 47.70 54.48 44.59 58.86
+Improvement +3.61 +8.68 +1.91 +4.65 +10.04 +9.72 +17.25 +20.53 +5.26 +6.41 +8.21 +9.99 +4.39 +7.94
+Backbone Designs Differences with Previous Study.
+Our results differ significantly from previous studies [35, 15], which were focused on LiDAR-based object detection tasks using CNN architecture with local receptive fields.
+In contrast, we focus on 3D SOT tasks within local search area and design a transformer architecture with global receptive fields.
+Therefore, by allocating more computational resources to the 1st stage, our modality-aware transformer-based backbone is able to capture the enough global-aware geometric semantics of the input point cloud.
+This is very different with CNN-based backbone, which require more computation overhead allocated to both the 1st, 2nd and 3rd stages.
+V Experiments
+Dataset and Evaluation metrics.
+Following previous works[12, 2, 4, 17, 22, 16].
+we evaluate our method on KITTI [41] and nuScenes [42] datasets.
+For KITTI datasets, we follow the same setting in baseline model[16] to divide the sequences into training, validation, and testing splits, where 0-16 for training, 17-18 for validation and 19-20 for testing.
+Notably, the nuScenes[42] is a more challenging dataset due to less channel (32-line) in LiDAR and more complex scenes, containing 1000 scenes, which are allocated into 700/150/150 scenes for training, validation, and testing, respectively.
+For metrics, we follow the baseline and use One-Pass Evaluation (OPE) [43] to measure Success and Precision.
+The Success measure the 3D IoU between the predicted box and the ground truth box, the Precision measures the AUC of the distance error between the center of the predicted and ground truth box.
+Implementation details.
+We follow our baseline model[16] and employ PVTv2-b2 [20] as backbone.
+The specific settings include a head number of [1, 2, 5, 8], a depth configuration of [3, 1, 1, 1], the expansion ratio for the feed-forward layer set to [8, 8, 4, 8] and a channel number configuration of [64, 128, 320, 512].
+In KITTI, we set [0.05, 0.05, 4] meters as pillar size and a search area of [-3.2, -3.2, -3, 3.2, 3.2, 1] meters for the car category.
+In nuScenes, we set [0.075, 0.075, 8] meters as pillar size and a search area of [-3.2, -3.2, -5, 3.2, 3.2, 3] meters for the car category.
+To fair comparison, other training setting is the same as our baseline model, more details refer to SMAT [16].
+V-A Comparison with State-of-the-art Trackers
+Quantitative results on KITTI.
+As shown in Table I, the proposed PillarTrack demonstrates significant performance improvements compared to the baseline method (SMAT [16]).
+Specifically, PillarTrack achieved a performance improvement of 11.4% 7.2% and 6.4% 4.2% in Success and Precision metrics in terms of mean by-class and mean-by frame setting, respectively.
+In the categories of the Car, Pedestrian, Van, and Cyclist, PillarTrack outperformed SMAT by 2.3%, 7.6%, 19.6%, and 16.8% in Success, respectively.
+Besides, PillarTrack also surpassed the motion-based method MMTrack [22] in all classes, achieving promising results on KITTI dataset.
+Quantitative results on NuScenes.
+The KITTI dataset has a limited number of samples, resulting in relatively modest performance gains, which may not fully reflect the capabilities of the model.
+Therefore, we also conduct experiments on the large-scale nuScenes [42] dataset.
+NuScenes contains many complex scenes with sparser point clouds, making it a more challenging benchmark.
+As shown in Table II, our PillarTrack achieves the second-best performance across the five categories on the nuScenes, demonstrating comparable performance among similarity-based methods.
+Notably, PillarTrack exhibits significant improvements over the baseline method [16] in each category, particularly with a 10.04% and 9.72% enhancement in Success and Precision metrics for the Truck category, and a 17.25% and 20.52% improvement for the Trailer category.
+However, our performance in the Car category is slightly lower than that of PTTR.
+We attribute this to PTTR’s utilization of both point-based and pillar-based representations for encoding the input point cloud, which leads to superior performance.
+Nevertheless, their point-based configuration may limit the potential for further quantization acceleration in onboard devices.
+Notably, M2-Tracker [22] adopts a motion-based paradigm for object tracking, which results in performance that surpasses similarity-based paradigms.
+TABLE III: Performance comparison of PE-PFE after noise disturbance.
+Baseline: original setting.
+Rotation: Add random rotation noise within -45° to 45°.
+Scale: Add random scaling from 0.8 to 1.2x for the input point clouds.
+Translation: Add random translation from -1.2m to 1.2m for the baseline.
+PFE Method Baseline Scale Translation Rotation
+Suc. Prec. Suc. Prec. Suc. Prec. Suc. Prec.
+PFE (SMAT[16]) 68.9 79.6 70.9 81.5 68.5 78.5 69.6 79.9
+PE-PFE (Ours) 70.8 80.3 72.9 82.4 73.3 83.0 72.0 81.5
+TABLE IV: Ablation of modality-aware transofrmer-based backbone. Bold denotes the best performance.
 Blocks Number GFLOPs Success Precision
 $\left[3,4,6,3\right]$ (SMAT[16]44footnotemark: 4) 0.37 68.9 79.6
-$\left[2,2,6,2\right]$ 0.29 25.7 34.9
-$\left[1,1,9,1\right]$ 0.29 53.1 56.2
+$\left[6,6,2,2\right]$ 0.38 70.2 80.5
 $\left[1,1,1,1\right]$ 0.14 69.5 80.1
@@
 $\left[1,1,1,4\right]$ 0.18 69.3 80.0
-To validate the aforementioned hypothesis for transformer-based backbone, we systematically study computation allocation by adjusting the stage compute ratio in the transformer-based backbone of SMAT [16].
-It is worth noting that the setting of SMAT has the same depth number as ResNet-50, with a stage compute ratio of $[3,4,6,3]$.
-As shown in Table II, we used the SMAT backbone as the baseline and varied the number of depth blocks from 0 to 4 with a stride of 1 in each stage.
-From the results, we can find that the performance is quite sensitive to the capacity of early stage 1, while being less sensitive to the later stages (2, 3, 4).
-This confirmed that allocating more computational resources to the early stages is beneficial for enabling the network to better capture the geometric information carried by the raw point clouds.
-Based on this observation, we set the stage compute ratio of the four stages to $[3,1,1,1]$.
-It is worth noting that with this depth blocks setting, our modality-aware transformer backbone network achieved promising performance improvements in terms of 3.2 point in Success and 2.5 point in Precision metric, while significantly reducing the computational cost by half the GFLOPs compared to the baseline method (from 0.37 to 0.18).
-Differences with Previous Study.
-Our results differ significantly from previous studies [44, 14], which were focused on LiDAR-based object detection tasks using CNN architecture with local receptive fields.
-In contrast, we focus on 3D SOT tasks within local search area and design a transformer architecture with global receptive fields.
-Therefore, by allocating more computational resources to the 1st stage, our network is able to capture the enough global-aware geometric information of the input point cloud.
-This is different with CNN-based backbone, which require more computational resources allocated to both the 1st and 2nd stages.
-As a result, our modality-aware transformer-based backbone design proves to be more effective.
-IV-C Activation Function Selection for Point Cloud
-[…]
-$ReLU(x)=\max(0,x)$ (2)
-$LeakyReLU(x)=\begin{cases}x&\text{if }x>0,\\ \alpha x&\text{if }x\leq 0.\end{cases}$ (3)
-There is a discrepancy between the choice of activation functions natural language processing (NLP) and vision architectures.
-Rectified Linear Unit (ReLU)[48] is extensively used in convolutional networks (ConvNets) due to its simplicity and efficiency.
-With the Transformer[26] merging, Gaussian Error Linear Unit (GELU)[49] is considered as a smoother variant of ReLU and has been employed in state-of-the-art Transformer models, such as Google’s BERT [50], OpenAI’s GPT-3 [51], and Vision Transformers [35, 34], demonstrating superior performance across various language and vision tasks.
-However, unlike image or text data, point cloud input data contains negative values that represent corresponding geometric property.
-We assume that directly applying ReLU or GELU activation functions to point clouds would suppress the negative values and result in information loss.
-To address this issue, we explored the Leaky Rectified Linear Unit (LeakyReLU)[52], which can address the issue of neuron death and preserve more information within the range of values less than zero.
-As shwon in Eq.
-1,2,3 The GELU and ReLU functions exhibit saturation behavior when the input values are less than zero, effectively reducing the information content.
-In contrast, the LeakyReLU function preserves the negative range of input values, enabling the retention of valuable information.
-TABLE III: Influence of different activation functions on Car category.Bold denotes the best result
-PFE Method Success Precision
-GELU (SMAT[16]44footnotemark: 4) 68.9 79.6
-ReLU 69.9 80.7
-LeakyReLU 71.2 (+2.3) 81.8 (+2.2)
-Figure 5: The impact of different activation functions on the reduction of the Epoch Loss.Inparticular, the LeakyReLU activation function has shown a more pronounced decrease in loss compared to the other two functions under consideration.
-Therefore, we simply replace the GeLU activatation function in SMAT with LeakyReLU function.
-As shown in Tab III and Figure 5, using LeakyReLU function has better performance (+2.3 in Success and 2.2 in Precision) and lower loss value than other functions, highlighting the significant impact of activation function choice on 3D single object tracking tasks based on point cloud modality.
-TABLE IV: Overall performance comparison on the KITTI dataset. The best two results are highlighted in red, blue.
-Method Publication Modality Paradigm View Mean-by class Mean-by frame
-Success Precision Success Precision
-$M^{2}$-Track  [21] CVPR 2022 LIDAR Motion Point 61.9 80.2 60.6 80.1
-SC3D[10] CVPR 2019 LIDAR Similarity Point 35.4 53.3 31.2 48.5
-P2B[2] CVPR 2020 Point 39.5 53.9 42.4 60.0
-3D-SiamRPN[1] IEEE Sensors J Point 43.8 58.6 46.6 64.9
-MLVSNet[53] ICCV2021 Point 44.1 60.3 45.7 66.7
-BAT[4] ICCV 2021 Point 47.2 65.1 51.2 72.8
-PTT[3] IROS 2021 Point 48.4 63.4 55.1 74.2
-LTTR[5] BMVC 2021 Voxel 50.0 67.4 48.7 65.8
-V2B[22] NeurIPS 2021 Point+Voxel 52.4 65.6 58.4 75.2
-PTTR[6] CVPR 2022 Point 58.4 77.8 57.9 78.1
-CMT[27] ECCV 2022 Point 57.2 76.0 59.4 77.6
-STNet[7] ECCV 2022 Point 63.4 81.4 61.3 80.1
-C2FT[54] RAL 2022 Point 51.6 67.7 57.2 76.4
-MLSET[55] RAL 2023 Point 54.2 68.9 59.6 78.4
-SMAT[16] RAL 2022 Pillar 56.7 76.1 60.4 79.5
-PillarTrack Ours Pillar 63.7 +7.0 78.9 +2.8 65.3 +4.9 82.2 +2.7
-TABLE V: Detailed per-class performance on the KITTI dataset. The best two results are highlighted in red, blue.
-Method Car-6424 Pedestrian-6088 Van-1248 Cyclist-308
-Success Precision Success Precision Success Precision Success Precision
-$M^{2}$-Track [21] 64.8 79.2 57.6 83.6 50.4 64.8 74.9 93.3
-SC3D [10] 41.3 57.9 18.2 37.8 40.4 47.0 41.5 70.4
-P2B [2] 56.2 72.8 28.7 49.6 40.8 48.4 32.1 44.7
-3D-SiamRPN[1] 58.2 76.2 35.2 56.2 45.7 52.9 36.2 49.0
-MLVSNet[53] 56.0 74.0 34.1 61.1 52.0 61.4 34.3 44.5
-BAT [4] 60.5 77.7 42.1 70.1 52.4 67.0 33.7 45.4
-PTT [3] 67.8 81.8 44.9 72.0 43.6 52.5 37.2 47.3
-LTTR [5] 65.0 77.1 33.2 56.8 35.8 45.6 66.2 89.9
-V2B [22] 70.5 81.3 48.3 73.5 50.1 58.0 40.8 49.7
-PTTR [6] 65.2 77.4 50.9 81.6 52.5 61.8 65.1 90.5
-CMT[27] 70.5 81.9 49.1 75.5 54.1 64.1 55.1 82.4
-STNet [7] 72.1 84.0 49.9 77.2 58.0 70.6 73.5 93.7
-C2FT [54] 67.0 80.4 48.6 75.6 53.4 66.1 38.0 48.7
-MLSET [55] 69.7 81.0 50.7 80.0 55.2 64.8 41.0 49.7
-SMAT [16] 71.9 82.4 52.1 81.5 41.4 53.2 61.2 87.3
-PillarTrack (Ours) 74.2 +2.3 85.1 +2.7 59.7+7.6 84.7 +3.2 43.4 +2.0 51.7 77.4 +16.2 94.2 +6.9
-V Experiments
-KITTI Dataset.
-We conduct training and evaluating for our task using the KITTI [56] datasets.
-We follow the same setting in baseline model[16] to divide the sequences into training, validation, and testing splits, where 0-16 for training, 17-18 for validation and 19-20 for testing.
-NuScenes Dataset.
-The NuScenes[57] dataset comprises 1000 scenes, which are allocated into 700/150/150 scenes for training, validation, and testing, respectively.
-Each scene is populated with approximately 300,000 points and is accompanied by comprehensive 360-degree viewpoint annotations.
-In our research, we adhere to the dataset configuration established by Open3DSOT[21] for the training and evaluation of our methodology.
-We also conduct a comparative analysis with the baseline model[16]to evaluate the efficacy and superiority of our designs.
-Evaluation metrics.
-We use One-Pass Evaluation (OPE) [58] to measure Success and Precision.
-The Success measure the 3D IoU between the predicted box and the ground truth box, the Precision measures the AUC of the distance error between the center of the predicted and ground truth box.
-Implementation details.
-We follow our baseline model[16] and employ PVTv2-b2 [37] as backbone, adhering to its original configuration.
-The specific settings include a head number of [1, 2, 5, 8], a depth configuration of [3, 1, 1, 1], the expansion ratio for the feed-forward layer set to [8, 8, 4, 8] and a channel number configuration of [64, 128, 320, 512].
-In KITTI dataset, we set [0.1, 0.1, 4] meters as pillar size and a search area of [-3.2, -3.2, -3, 3.2, 3.2, 1] meters for the car category.
-In nuScenes dataset, we set [0.075, 0.075, 8] meters as pillar size and a search area of [-3.2, -3.2, -5, 3.2, 3.2, 3] meters for the car category.
-Additionally, the LeakyReLU activation is employed with a negative slope value of 0.01.
-Training details.
-On the KITTI dataset, we train PillarTrack with AdamW[59] optimizer with an initial learning rate of 0.0001, a weight decay of 0.05, and a batch size of 16 on a NVIDIA 3090 GPU.
-The total number of training epochs was set to 72.
-For the NuScenes dataset, the training was conducted over 18 epochs, with a learning rate decay by a factor of 10 at epoch 14.
-Method Paradigm Car-64159 Pedestrian-33227 Truck-13587 Trailer-3352 Bus-2953 Mean by class Mean by frame
-Success Precision Success Precision Success Precision Success Precision Success Precision Success Precision Success Precision
-$M^{2}$-Tracker [21] Motion 55.85 65.09 32.10 60.92 57.36 59.54 57.61 58.26 51.39 51.44 50.86 59.05 49.23 62.73
-SC3D [10] Similarity 22.31 21.93 11.29 12.65 30.67 27.73 35.28 28.12 29.35 24.08 25.78 22.90 20.70 20.20
-P2B [2] 38.81 43.18 28.39 52.24 42.95 41.59 48.96 40.05 32.95 27.41 38.41 40.90 36.48 45.08
-PTT [12] 41.22 45.26 19.33 32.03 50.23 48.56 51.70 46.50 39.40 36.70 40.38 41.81 36.33 41.72
-BAT [4] 40.73 43.29 28.83 53.32 45.34 42.58 52.59 44.89 35.44 28.01 40.59 42.42 38.10 45.71
-GLT-T [60] 48.52 54.29 31.74 56.49 52.74 51.43 57.60 52.01 44.55 40.69 47.03 50.98 44.42 54.33
-PTTR [6] 51.89 58.61 29.90 45.09 45.30 44.74 45.87 38.36 43.14 37.74 43.22 44.91 44.50 52.07
-SMAT [16] 43.51 49.04 32.27 60.28 44.78 44.69 37.45 34.10 39.42 34.32 39.49 44.49 40.20 50.92
-PillarTrack 47.12 57.72 34.18 64.93 54.82 54.41 57.70 54.63 44.68 40.73 47.70 54.48 44.59 58.86
-Improvement +3.61 +8.68 +1.91 +4.65 +10.04 +9.72 +17.25 +20.53 +5.26 +6.41 +8.21 +9.99 +4.39 +7.94
-TABLE VI: Performance comparison on the NuScenes dataset. The best two results are highlighted in red, blue.
-V-A Quantitative results on KITTI.
-As shown in Table IV and V, the proposed PillarTrack demonstrates significant performance improvements compared to the baseline method (SMAT [16]).
-Specifically, PillarTrack achieved a performance improvement of 7.0% 2.8% and 4.9% 2.7% in Success and Precision metrics in terms of mean by-class and mean-by frame setting, respectively.
-In the categories of the Car, Pedestrian, Van, and Cyclist, PillarTrack outperformed SMAT by 2.3%, 7.6%, 2.0%, and 16.2% in Success, respectively.
-These improvements can be attributed to the effective designs of PillarTrack.
-Firstly, PE-PFE module could stabilize network training without information loss.
-Secondly, the modality-aware transformer-basde backbone focuses on capturing more geometric information from point clouds and enhance the feature representation capability of pillars.
-Lastly, LeakyReLU activation retains more information and help network optimization.
-Besides, PillarTrack also surpassed the motion-based method MMTrack [21] in the categories of Car, Pedestrian, and Cyclist, achieving state-of-the-art performance on the KITTI dataset.
-However, for Van, our method only has 43.4% in Success.
-We believe there are two main reasons for this.
-Firstly, the point clouds of vans are concentrated on the object surfaces, and after pillar encoding, they lack distinctive geometric features, making them more challenging to track accurately.
-Secondly, the limited number of training samples for the Van category may have constrained the model’s performance, as it might not have captured the full range of variations and characteristics exhibited by vans in the real world.
-V-B Quantitative results on NuScenes.
-As illustrated in Table VI, our PillarTrack achieves the second-best performance across the five categories on the nuScenes dataset.
-Compared to previous methods, M2-Tracker[21] adopts a different motion-based paradigm for object tracking, demonstrating superior results compared to the similarity-based paradigm.
-Meanwhile, our PillarTrack attains the best results among the similarity-based methods.
-Furthermore, compared to the baseline method[16], PillarTrack has demonstrated significant improvements in each category.
-Specifically, PillarTrack has enhanced the mean class-wise and frame-wise Success and Precision metrics by 8.21% / 9.99% and 4.39% / 7.94%, respectively.
-In terms of Success for each category, PillarTrack achieves notable improvements of 3.61% for Car, 1.91% for Pedestrian, 10.04% for Truck, 17.25% for Trailer, and 5.26% for Bus.
-Importantly, the Truck and Trailer categories demonstrate significantly enhanced performance compared to baseline.
-However, such improvements were not observed in the Van category in the KITTI dataset.
-We attribute this to two reasons: (1) The KITTI dataset only provides front-view annotations.
-Under the front-view perspective, the model is constrained to learn feature representations specific to the truck from a single viewpoint, which limits the acquisition of comprehensive geometric information about the Van.
-As a result, the Siamese network based on appearance matching struggles to capture the key geometric information of the Van.
-Conversely, the NuScenes dataset provides 360-degree surround view point cloud data with corresponding annotations.
-This enables superior learning of both appearance and geometric representations, leading to significant performance enhancement by effectively utilizing multi-view point cloud data.
-(2) The KITTI dataset contains a limited number of annotated van samples (1248), whereas the NuScenes dataset offers a larger sample size (13587/3352).
-Within our effective framework design, this facilitates richer feature representation learning, leading to performance improvements.
-Besides, our performance is slightly lower than that of PTTR on Car category.
-We think this is because PTTR utilizes both point-based and pillar-based representations to encode the input point cloud, resulting in superior performance, particularly in terms of the Success metric.
+V-B Ablation Study
+The robustness of Pe-PFE.
+As shown in Tab.
+III, we evaluated the performance of our PE-PFE under different noise disturbance, with a particular focus on its robustness against scale, translation, and rotation variations.
+The results demonstrate that PE-PFE outperforms the baseline method PFE (SMAT[16]) across all conditions.
+These results indicate that PE-PFE effectively maintains performance stability in the presence of various noise disturbances, showcasing strong robustness.
+The Design of Modality-aware Transformer-based Backbone.
+We conduct experiments by arranging the number of blocks per stages from 0 to 4 with a stride of 1, while setting the number of blocks in other stages to 1, as shown in Tab.
+IV.
+The results prove our previous assumption: for pillar-based LiDAR 3D SOT tasks, we should reallocate the capacity to the early stages to better integrate the geometric information carried by the raw points, instead of allocating the capacity in the later stages like image domain.
+Therefore, to achieve a better performance under the constraint of low computation resource, we set the stage compute ratio to (3, 1, 1, 1).
+TABLE V: Ablation experiment with Blocks Number and PFE Method in the Car category in KITTI.
+Method Blocks Number PFE Method Success Precision
+SMAT[16] $\left[3,4,6,3\right]$ PFE 68.9 79.6
+Ours $\left[3,4,6,3\right]$ PEPFE 70.8 80.3
+$\left[3,1,1,1\right]$ PFE 72.1 82.1
+$\left[3,1,1,1\right]$ PEPFE 74.2 85.1
+Overall Ablation.
+Here, we conducted an ablation study to investigate the impact of the proposed PE-PFE and modality-aware backbone network design on the overall performance, as shown in Table V.
+Firstly, when employing the same block configuration ($\left[3,4,6,3\right]$), the PE-PFE method achieved a success rate of 70.8% and a precision of 80.3%, surpassing the baseline method SMAT.
+Secondly, by adjusting the block configuration to $\left[3,1,1,1\right]$, we observed that the success rate of the PE-PFE method further increased to 74.2%, with precision rising to 85.1%.
+This indicates that reducing the number of blocks, in conjunction with the PE-PFE strategy, effectively enhances the model’s tracking capability.
+Figure 6: Visualization results comparison of our PillarTrack and baseline method (SMAT[16]) in KITTI dataset.
 V-C Visualization results.
-Figure 6: Visualization results comparison of our PillarTrack and baseline method (SMAT[16]) on the categories of Car, Pedestrian and Cyclist in KITTI dataset.
-To comprehensively evaluate the performance of our method, particularly its capability to handle point clouds of varying sparsity, we conduct visualization on the different category, encompassing both sparse and dense scenes.
-Through comparative experiments on the initial frame and subsequent sequences, our method demonstrated superior tracking accuracy in both sparse and dense point cloud environments compared to the baseline SMAT method, highlighting ours enhanced adaptability and robustness.
-In the Cyclist category, our method was able to consistently track the target, whereas SMAT occasionally lost the target.
-For the Pedestrian category, our method maintained stable tracking from the outset, while SMAT failed to accurately identify the target initially.
-The visualization results in Figure 5 further substantiate the effectiveness and superior performance of our method across various scenarios.
-V-D Running Speed
+To comprehensively evaluate the performance of our method, particularly its more representative capability of point clouds, we conduct visualization on the different category, encompassing both sparse and dense scenes.
+Through comparative experiments on the initial frame and subsequent sequences, our method demonstrated superior tracking accuracy in both sparse and dense scenes compared to the baseline[16] method, highlighting ours enhanced adaptability and robustness.
+In the Cyclist category, our method was able to consistently track the target, whereas SMAT lost the target.
+For the Pedestrian category, our method maintained stable tracking from the outset, while SMAT failed to identify the targets.
+The visualization results in Figure 6 substantiate the effectiveness and superior performance of our method across various scenarios.
+V-D Running Speed.
 We calculated the average running time of all test frames for car to measure PillarTrack’s speed.
-PillarTrack model achieved 27 FPS on a single NVIDIA 3080 GPU, while baseline model SMAT has a 18.2 FPS in the inference phase.
+PillarTrack model achieved 27 FPS on a single NVIDIA 3080 GPU, while baseline model SMAT only has 18.2 FPS.
 We did not introduce extra computational parameters, and we further reduced the computational overhead on the backbone while increasing the model’s performance.
 This once again demonstrates the effectiveness of our PillarTrack.
-Furthermore, our proposed PillarTrack can leverage quantization techniques[61, 62] to accelerate the runtime.
-We leave this problem for future research.
+Furthermore, our PillarTrack can leverage quantization techniques[44] to accelerate the runtime.
+We leave this problem for the future research.
 VI CONCLUSIONS
 In this work, we proposed PillarTrack, a pillar-based 3D single-object tracking framework that decreases computation overhead while improving performance.
-The framework introduces a Pyramid-type Encoding Pillar Feature Encoder (PE-PFE) and a modality-aware Transformer-based backbone, and a simply replacement of activation function.
-Our extensive experiments on the KITTI dataset demonstrate PillarTrack’s superior performance, achieving an better trade-off between speed and accuracy.
-Based on above designs and its efficiency, We hope to inspire the community to rethink the design of network for 3D single object tracking based on point clouds.
+We propose a Pyramid-Encoding Pillar Feature Encoder (PE-PFE) and a modality-aware Transformer-based backbone to enhance the pillar representation.
+Extensive experiments on the KITTI and nuScenes demonstrate the superior performance of PillarTrack.
+Based on above designs and its efficiency, We hope our method could provide new insights for the design of 3D SOT network based on point clouds.